# Auditing for Racial Discrimination in the Delivery of Education Ads

### v1

@@ 1. Introduction
Why Education? In this work, we explore education as a new domain for which to study the potential role of ad delivery for discrimination. Avoiding discrimination in the delivery of educational opportunities matters because education has long-term impact on the career and financial well-being of individuals (Federal Reserve, 2023). One specific concern in U.S. higher education is the potential for for-profit colleges to produce poor financial outcomes for their graduates (Looney and Yannelis, 2015). Such colleges often devote considerable resources to advertising, targeting racial minorities disproportionately (Kahn et al., 2019). Prior work has not studied whether ad delivery algorithms propagate historical racial targeting biases in such colleges, a unique aspect of the education domain. In addition, education is an important category because of advertiser large spending in this domain. According to Meta’s whitepaper ((Wernerfelt et al., 2022), Figure 4(a)), education is one of the top eight verticals among its advertisers, significantly ahead of verticals such as political advertising which have received much scrutiny.
@@
Finally, we use our methodology to investigate whether Meta’s algorithm steers delivery of ads for schools with historical predatory marketing practices disproportionately towards Black individuals. Over-delivery of ads for such schools could disproportionately harm students who are unaware of their documented risks to attendees (Kahn et al., 2019; Deming et al., 2013). We pick ads for three for-profit schools that were subject to legal action by the U.S. Department of Education due to their predatory practices (of Education, 2018), and evaluate the racial difference in delivery of their ads compared to delivery of ads for public schools. We find that Meta shows the ads for the schools with historical predatory practices to relatively more Black individuals. Our results show that it is not sufficient for schools to target their ads equitably, but that ad platforms also need to ensure their algorithms are not introducing biases along legally protected characteristics such as race.
@@
2. Methodology: Using Pair of Schools with De-Facto Racial Skew in Recruitment
@@ 2.3. Running Ads on Meta and Evaluating their Delivery
Having described the key insight of our methodology based on which we identify pairs of education opportunities to advertiser for, we next describe the steps we take to actually run the ads on Meta’s ad platform.
@@ 2.3.2. Selecting Ad Creatives and Campaign Parameters
Realistic creatives: To test the effects of ad delivery for real-world ads used by schools, we also run experiments using realistic ad creatives. We take a snapshot of a list of ads run by the schools from Meta’s public Ad Library (Meta, 2024b) and manually annotate the images used in the ads by (perceived) race. The manual effort introduces the potential for annotation bias but extreme precision is not mandatory for our methodology. For each school, we select from the list one representative ad that includes the face of an individual whose race is represented in majority of the ads. Figure 2(c) and Figure 2(d) show an example pair of realistic ad creatives we use in our experiments.
@@ 3.2. Demonstrating Amplified Skew Using Realistic Ad Creatives
3.3. Experiments Using Schools with Historical Predatory Practices
We next apply our methodology to examine whether Meta’s algorithm delivers ads for for-profit schools with historical predatory marketing practices to disproportionate fraction of Black individuals. Disproportionate promotion of opportunities at such institutions poses potential harm to students who enroll in these schools (Deming et al., 2013; Looney and Yannelis, 2015; Kahn et al., 2019).
We find historically predatory schools based on a list published by the U.S. Department of Education of schools that previously have been sued or fined for deceptive marketing practices (of Education, 2018). From this list, we pick the three universities with the largest number of students: DeVry, Grand Canyon and Keiser. We then evaluate ad delivery for those schools using our methodology. We would like to test whether Meta’s algorithms perpetuate the historical skew even if the schools have improved their marketing practices in response to the legal challenges.
Table 3. Racial make-up of historically predatory for-profit schools and the public schools we pair them with.
Pair ID For-profit (historically predatory) school Public school
@@ 3.3. Experiments Using Schools with Historical Predatory Practices
We apply the methodology described in § 2, but we modify the school-selection criteria to consider historically predatory schools instead of those with skew in racial demographics. We continue to match these for-profit schools with the same three public schools used in our earlier experiments, following the process outlined in § 2.2. Table 3 provides an overview of the school pairings we used in this experiment. Similar to § 3.2, we use realistic ad creatives that are taken from each school’s Meta ad library page (see Appendix B for ad screenshots). As in our previous experiments, we replicate the experiments on a “flipped” audience to ensure location does not skew our results.
We find that, for all three pairs, the historically predatory for-profit school ad is shown to relatively larger fraction of Blacks compared to the public school ad (Figure 5; left column). For example, for the first pair of schools (top row of Figure 5(a)), the public school ad (orange circle) is delivered to 44% Blacks whereas the for-profit school ad (blue cross mark) is delivered to relatively larger fraction of Blacks (52%). For two out of the three pairs, the racial skew in delivery is statistically significant (see Figure 5(d)) and remains statistically significant after applying Holm’s correction.
In summary, our findings show that Meta’s algorithms disproportionately deliver educational opportunities that could be harmful to individuals based on race. Our study illustrates how this harm can happen in practice, as the for-profit schools with historical predatory advertising are currently active advertisers on Meta’s ad platform.
@@
Figure 5. Skew measured by comparing the delivery of ads for a historically predatory for-profit school and a public school.
@@ 4.1. Platforms Shaping Access to Education Opportunities
Our findings highlight the negative impact of ad delivery algorithms in shaping access to education opportunities, adding to existing concerns about the disadvantages of for-profit schools and their historical involvement in predatory advertising. Studies have raised concerns that for-profit colleges provide poorer outcomes, with their students showing higher loan default rate, lower earnings and employment than comparable students at other post-secondary institutions (Deming et al., 2013; Looney and Yannelis, 2015; Federal Reserve, 2023). In addition, it is known that for-profit colleges have historically engaged in predatory marketing that targeted racial minorities (Kahn et al., 2019), and enroll a disproportionate number of students with low family income (Chetty et al., 2017).
The platform-induced racial bias we demonstrate illustrates another significant factor that decides exposure to education opportunities: ad delivery algorithms. We show Meta’s algorithms deliver ads for the for-profit schools to relatively more Black individuals than ads for public schools. The racial difference we observe in the delivery is not due to the advertiser’s targeting choices since we select racially balanced audiences. It is also not due to market effects or difference in platform use by race since our methodology controls for those. Therefore, even if for-profit institutions aim for racially balanced ad targeting, Meta’s algorithms would recreate historical racial skew in who the ad are shown to, and would do so unbeknownst to the advertisers. Our findings show it is not enough for schools to target their ads equitably; platforms also need to ensure their ad delivery algorithms are not biased by race.
@@ Appendix A Experiments Using State Rather than DMA when Constructing Audiences
Similar to our findings in § 3.2 and § 3.1, these experiments show evidence of racial bias in the delivery of education ads (left column of Figure 6), and show the degree of bias in delivery increases when we use realistic ad creatives (right column of Figure 6). However, we see some variation in the experiments we replicate on flipped audiences for which we expected the results to be similar. For example, in Figure 6e, approximately 52% of recipients of the for-profit school ad are Black for aud5, whereas 70% or recipients are Black when the audience is flipped (aud5f). Because we do not look at absolute numbers but rather relative differences in the delivery of the for-profit and public schools, these variations do not affects our evaluation of skew in the algorithm. However, the results show using states when constructing audiences can introduce additional variations one must keep into account.
Appendix B Realistic Ad Creatives Used for Historically Predatory For-profit Schools
In Figure 7, we show the realistic ad creatives used for experiments involving historically predatory for-profit schools in § 3.3. These ad creatives are taken from each school’s page on Meta’s public ad library (see § 2.3.2 for details on creative choice).
@@ Appendix B Realistic Ad Creatives Used for Historically Predatory For-profit Schools
Figure 7. Realistic ad creatives used for for-profit schools with historical practices of predatory marketing. The ad texts and images are taken from each school’s page on Meta’s ad library.

### v2

@@ 1. Introduction
Why Education? In this work, we explore education as a new domain for which to study the potential role of ad delivery for discrimination. Avoiding discrimination in the delivery of educational opportunities matters because education has long-term impact on the career and financial well-being of individuals (Federal Reserve, 2023). One specific concern in U.S. higher education is the potential for for-profit colleges to produce poor financial outcomes for their graduates (Deming et al., 2013; Looney and Yannelis, 2015). Such colleges spend a large amount on advertising (Deming et al., 2013), and have been accused of disproportionately targeting racial minorities (Kahn et al., 2019). Prior work has not studied whether ad delivery algorithms propagate historical racial targeting biases in such colleges, a unique aspect of the education domain. In addition, education is an important category because of advertiser large spending in this domain. According to Meta’s whitepaper ((Wernerfelt et al., 2022), Figure 4(a)), education is one of the top eight verticals among its advertisers, significantly ahead of verticals such as political advertising which have received much scrutiny.
@@
Finally, we use our methodology to investigate whether Meta’s algorithm steers delivery of ads disproportionately by race for schools whose enrollment or marketing practices have previously received legal scrutiny. Over-delivery of ads for such schools may inadvertently disadvantage students who are not aware of the potential risks, such as poor labor market outcomes (Deming et al., 2013). We pick ads for three such schools from a list compiled by Veterans Education Success in 2018 (Success, 2018) and evaluate the racial difference in delivery of their ads compared to delivery of ads for public schools. We find that Meta shows the ads for the selected schools to relatively more Black individuals. Our results show that it may not be sufficient for these schools to target their ads equitably, but that ad platforms also need to ensure their algorithms are not introducing biases along legally protected characteristics such as race.
@@
2. Methodology: Using Pair of Schools with De-Facto Racial Skew in Enrollment
@@ 2.3. Running Ads on Meta and Evaluating their Delivery
Having described the key insight of our methodology based on which we identify pairs of education opportunities to advertise for, we next describe the steps we take to actually run the ads on Meta’s ad platform.
@@ 2.3.2. Selecting Ad Creatives and Campaign Parameters
Realistic creatives: To test the effects of ad delivery for real-world ads used by schools, we also run experiments using realistic ad creatives. We take a snapshot of a list of ads run by the schools from Meta’s public Ad Library (Meta, 2024b) and manually annotate the images used in the ads by (perceived) race. The manual effort introduces the potential for annotation bias but extreme precision is not mandatory for our methodology. For each school, we select from the list one representative ad that includes the face of an individual whose race is represented in majority of the ads. 2(c) and 2(d) show an example pair of realistic ad creatives we use in our experiments.
@@ 3.2. Demonstrating Amplified Skew Using Realistic Ad Creatives
3.3. Experiments Using Schools with Prior Legal Scrutiny
We next apply our methodology to examine whether Meta’s algorithm racially biases the delivery of ads for schools with prior legal scrutiny for their marketing or recruitment practices. Prior work has suggested that opportunities at such institutions can provide poor financial outcomes to students who enroll in these schools (Deming et al., 2013; Looney and Yannelis, 2015).
We identify schools that have previously received legal scrutiny for their marketing and recruitment practices from a list published by Veterans Education Success in 2018 for the U.S. Department of Education National Advisory Committee on Institutional Quality and Integrity (Success, 2018). From this list, we pick three private universities with the largest number of students: DeVry, Grand Canyon and Keiser. DeVry and Grand Canyon are currently private for-profit universities; Keiser was a private, for-profit university previously but changed to non-profit status in 2011. We then evaluate delivery of ads advertising educational opportunities at those schools using our methodology. Our goal is to test whether Meta’s algorithms result in skewed delivery even if the schools have changed their marketing practices.
Table 3. Racial make-up of private schools with prior legal scrutiny and the public schools we pair them with.
Pair ID Private School Public School
@@ 3.3. Experiments Using Schools with Prior Legal Scrutiny
We apply the methodology described in § 2, but we modify the school-selection criteria to consider the three schools we selected as described above. We continue to match the selected schools with the same three public schools used in our earlier experiments, following the process outlined in § 2.2. Table 3 provides an overview of the school pairings we used in this experiment. Similar to § 3.2, we use realistic ad creatives that are taken from each school’s Meta ad library page (see Appendix B for ad screenshots). As in our previous experiments, we replicate the experiments on a “flipped” audience to ensure location does not skew our results.
We find that, for all three pairs, the ad for the school with prior legal scrutiny is shown to relatively larger fraction of Blacks compared to the public school ad (Figure 5; left column). For example, for the first pair of schools (top row of Figure 5(a)), the public school ad (orange circle) is delivered to 44% Blacks whereas the ad for the private school (blue cross mark) is delivered to relatively larger fraction of Blacks (52%). For two out of the three pairs, the racial skew in delivery is statistically significant (see Figure 5(d)) and remains statistically significant after applying Holm’s correction.
Our findings show that Meta’s algorithms disproportionately deliver educational opportunities at the selected schools based on race. Our study illustrates how platform-driven discrimination can happen in practice, as the schools are active advertisers on Meta’s ad platform.
@@
Figure 5. Skew measured by comparing the delivery of ads for private schools with prior legal scrutiny and public schools.
@@ 4.1. Platforms Shaping Access to Education Opportunities
Our findings highlight the negative impact of ad delivery algorithms in shaping access to education opportunities, adding to existing concerns about the disadvantages of attending for-profit schools. Studies have raised concerns that for-profit colleges provide poorer outcomes, with their students showing higher loan default rate, lower earnings and employment than comparable students at other post-secondary institutions (Deming et al., 2013; Looney and Yannelis, 2015; Federal Reserve, 2023). Kahn et al. have highlighted the risk that for-profit schools result in high debt and negative outcomes, particularly for Black students (Kahn et al., 2019).
The platform-induced racial bias we demonstrate illustrates another significant factor that decides exposure to education opportunities: ad delivery algorithms. We show Meta’s algorithms deliver ads for the for-profit schools to relatively more Black individuals than ads for public schools. The racial difference we observe in the delivery is not due to the advertiser’s targeting choices since we select racially balanced audiences. It is also not due to market effects or difference in platform use by race since our methodology controls for those. Therefore, even if an institution with past skewed racial enrollment aims for racially balanced ad targeting, Meta’s algorithms would recreate historical racial skew in who the ad are shown to, and would do so unbeknownst to the advertisers. Our findings show it is not enough for schools to target their ads equitably; platforms also need to ensure their ad delivery algorithms are not biased by race.
@@ Acknowledgements.
Errata
The arXiv v2 version of this paper was updated from arXiv v1 to clarify language related to experiments in § 3.3.
@@ Appendix A Experiments Using State Rather than DMA when Constructing Audiences
Similar to our findings in § 3.2 and § 3.1, these experiments show evidence of racial bias in the delivery of education ads (left column of Figure 6), and show the degree of bias in delivery increases when we use realistic ad creatives (right column of Figure 6). However, we see some variation in the experiments we replicate on flipped audiences for which we expected the results to be similar. For example, in Figure 6(e), approximately 52% of recipients of the for-profit school ad are Black for aud5, whereas 70% or recipients are Black when the audience is flipped (aud5f). Because we do not look at absolute numbers but rather relative differences in the delivery of the for-profit and public schools, these variations do not affects our evaluation of skew in the algorithm. However, the results show using states when constructing audiences can introduce additional variations one must keep into account.
Appendix B Realistic Ad Creatives Used for Private Schools with Prior Legal Scrutiny
In Figure 7, we show the realistic ad creatives used for experiments involving the private schools selected in § 3.3. These ad creatives are taken from each school’s page on Meta’s public ad library (see § 2.3.2 for details on creative choice).
@@ Appendix B Realistic Ad Creatives Used for Private Schools with Prior Legal Scrutiny
Figure 7. Realistic ad creatives used for private schools selected in § 3.3. The ad texts and images are taken from each school’s page on Meta’s ad library.